OpenStudy (anonymous):

Given the sequence 2, 8, 32, 128, 512 … , what is S10 ?

OpenStudy (anonymous):

524,288

OpenStudy (anonymous):

That's what i got but its not in the multiple choice answers.. :( A. 174,762 B. 699,050 C. 2,796,202 D. 1,073,741,823

OpenStudy (zzr0ck3r):

\(\sum_{n=1}^{10}2^{2n-1}=699050\)

OpenStudy (anonymous):

Well i guess this guy is right.

OpenStudy (anonymous):

So its B... Im still confused on how though to be honest.

OpenStudy (zzr0ck3r):

do you get why I used \(2^{2n-1}\)?

OpenStudy (anonymous):

Yeah.. i think.

OpenStudy (zzr0ck3r):

well then, what is confusing? I just added them up

OpenStudy (anonymous):

Okay now i get it. I guess I was confused on how you found that in the first place.... but i get it now. Thanks!

OpenStudy (zzr0ck3r):

\(2^{2*1-1}+2^{2*2-1}+.....+2^{2*10-1}\)

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

you just have to see the pattern we have 2 = 2^1 8=2^3 32=2^5 ..... so we needed something that would index 1,3,5.... and 2n-1 will do that.